1. In class, we saw that for a fall in price the Compensating Variation (CV) was less than Marshallian Consumer Surplus (MCS), and that MCS was less than Equivalent Variation (EV). Use a 2-panel graph to demonstrate that CV>MCS>EV for a price increase.
Answer #1

CV < MCS < EV. This relationship is true when the good is a normal good and the price of the good falls.

In case of price increase, following relationship holds

CV > MCS > EV

For a normal good, the hicksian demand curve is usually more steeper than the Marshallian demand curve. This is shown in the below figure.

An increase in the price shifts the demand curve to the left.

The area of CV due to this price increase = p+q+r

CV is obtained at the new price level with the old level of utility

MCS = p+q

For Marshallian Demand, the income is fixed at I bar.

EV = p

It is obtained at new utility with the old price level.

So, p+q+r > p+q > p

CV > MCS > EV

3470 (us Pipe) p.) ... q=D (I, Piipe) na x . ă ţ uo = Old utility O= New utility due to price increase